OpenStudy (anonymous):

Ben has a 25 foot ladder leaning on the side of his house. If the base of the ladder is 7 feet from the house, how far off the ground is the top of the ladder? Solve.

OpenStudy (anonymous):

7^2+ b^2= 25^2 49+b^2 =625 625- 49= your answer

OpenStudy (anonymous):

sqrt(25^2 - 7^2) (Pythagoras)

OpenStudy (anonymous):

you know the length of 2 of the sides, use pythagorean theory to figure out the third side, its like a triangle